OpenStudy (anonymous):

If f(x) = integral of 1/ (sqrt of t^3 +2)dt, then f'1 = ? I'll write the equation with teh equation thing also

OpenStudy (turingtest):

is the 2 under the radical sign?

OpenStudy (anonymous):

\[If f(x) = \int\limits_{0}^{2x} 1/ (\sqrt{t^3+2}) dt, \] then f' (1) = ?

OpenStudy (turingtest):

oh it's a fundamental theorem of calculus problem right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

the upper limit says 2x btw

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

help me!

OpenStudy (turingtest):

Well this problem is gonna bug me if I don't figure it out, so give me a sec...

OpenStudy (anonymous):

r u still there

OpenStudy (turingtest):

yes, but I'm a bit thrown by how they want you to get the value of the integral... perhaps I'm just out of practice on these :/

OpenStudy (anonymous):

=( so u dont know how 2 approach teh problem?

OpenStudy (anonymous):

ok the derivative is the integrand, but you need the chain rule here

OpenStudy (anonymous):

\[ f(x) = \int\limits_{0}^{2x} \frac{1}{\sqrt{t^3+2}}dt,\] \[f'(x)=\frac{1}{\sqrt{(2x)^3+2}}\times 2\] via the chain rule. clean this up with some algebra or just replace x by 1 to get \[f'(1)\]

OpenStudy (turingtest):

oh right, so the f(0) doesn't matter then...

OpenStudy (anonymous):

i get \[f'(1)=\frac{2}{\sqrt{10}}\]

OpenStudy (turingtest):

f'(0) I mean*

OpenStudy (anonymous):

\[f(0)=0\]

OpenStudy (turingtest):

? how do you know?

OpenStudy (anonymous):

the lower limit of integration gives you the constant. the only work here to to write the derivative making sure to use the chain rule because the upper limit is 2x and not x

OpenStudy (turingtest):

oh right XD

OpenStudy (anonymous):

\[f(0)=\int_0^0f(t)dt=0\]

OpenStudy (turingtest):

so it's not that f(0) is zero by itself, it disappears because we took the derivative and the derivative of a constant is zero

OpenStudy (anonymous):

\[f(x)=\int_a^xf(t)dt\] is a function of x. if you replace x by a you get \[f(a)=\int_a^af(t)dt=0\]

OpenStudy (turingtest):

ok I see what you're saying...

OpenStudy (anonymous):

that is how we know that for example \[\int_1^2t^3dt=\frac{2^4}{5}-\frac{1}{5}\] because we know that both \[f(x)=\int_1^2t^3dt\] and \[g(x)=\frac{x^3}{4}+C\] have the same derivative, and we also know that \[f(1)=0\] meaning that \[g(1)=0=\frac{1}{4}+C\]

OpenStudy (anonymous):

typo up top, denominators should be 4!

OpenStudy (turingtest):

f(x) can't have constant bounds on both ends though can it? that makes it a constant, 7/2, the value of the integral

OpenStudy (turingtest):

so the derivative of f(x) the way you wrote it would be zero ...no? sorry to carry on in your post sabrina

OpenStudy (anonymous):

not sure exactly what you mean \[\int_a^b f(t)dt\] is a number, whereas \[\int_a^xf(t)dt\] is a function of x

OpenStudy (anonymous):

and the derivative of \[\int_a^xf(t)dt\] is \[f(x)\] by ft of c

OpenStudy (anonymous):

no problem..u asking questions calrifies many things for me also

OpenStudy (anonymous):

maybe i can explain this way, see if it makes sense. with a simple example. suppose i write \[f(x)=\int_1^x2tdt\] this is a function of x. and clearly \[f(1)=\int_1^12tdt=0\] now suppose i would like to find \[f(3)=\int_1^32tdt\] i know two functions whose derivative is \[2x\] one is \[f(x)=\int_1^x2tdt\] and the other is \[x^2\]

OpenStudy (saifoo.khan):

sat when u are free, please see my last question

OpenStudy (anonymous):

because i have two functions with the same derivative, i know they can at most differ by a constant. and what is that constant? well since \[f(1)=0\] i know the constant must be \[-1\] since if i evaluate \[x^2\] at 1 i get 1 and thefore i know that \[f(x)=\int_1^x2tdt=x^2-1\]

OpenStudy (anonymous):

this allows me to evaluate \[f(x)\] at any number i choose. so \[f(3)=\int_1^32tdt=3^2-1\]and so on

OpenStudy (turingtest):

but you wrote above\[f(x)=\int_{1}^{2}t^3dt\]which is a constant I think it was just a typo, I understand the rest of what you are saying thanks I really do get it better now.